### v2

# Autonomous Learning with High-Dimensional Computing Architecture Similar to von Neumann’s

Abstract
We model human and animal learning by computing with high-dimensional vectors ($D$ = 10,000 for example). The architecture resembles traditional (von Neumann) computing with numbers, but the instructions refer to vectors and operate on them in superposition. The architecture includes a high-capacity memory for vectors, counterpart of the random-access memory (RAM) for numbers. The model’s ability to learn from data reminds us of deep learning, but with an architecture closer to biology. The architecture agrees with an idea from psychology that human memory and learning involve a short-term working memory and a long-term data store. Neuroscience provides us with a model of the long-term memory, namely, the cortex of the cerebellum. With roots in psychology, biology, and traditional computing, a theory of computing with vectors can help us understand how brains compute. Application to learning by robots seems inevitable, but there is likely to be more, including language. Ultimately we want to compute with no more material and energy than used by brains. To that end, we need a mathematical theory that agrees with psychology and biology, and is suitable for nanotechnology. We also need to exercise the theory in large-scale experiments. The analogy with traditional computing suggests that the architecture be programmable in terms of variables, values and data structures, the very things that have made traditional computing ubiquitous and that seem worth learning from and emulating.
1 Preamble
It seems best to clear a likely misconception up front. We generally agree that the everyday computer is a “von Neumann machine,” and we call other architectures “non-von-Neumann,” including Artificial Neural Nets (ANNs). Computing with wide vectors, as discussed in this paper, grew out of artificial neural nets and is an ANN of a kind, but it realizes a von Neumann-like architecture with an instruction set and a memory for wide vectors, and it allows variables and values to be used in traditional ways. Three vector operations are used to encode and decode composite structure, and to capture probabilities in bit counts and vector similarity. Minimal reliance on backpropagation is another thing that sets the “new” architecture apart from mainstream ANNs. In abstract theory terms, the architecture is no less powerful than an ordinary computer, or the Universal Turing Machine without an endless tape. The finite automaton at the core is simple and easily realized in an algebra of vectors. The point is to program with vectors as basic objects, much as we now program with numbers.
2 Background
We explore the idea that human beings, animals, and robots can learn to act intelligently in the world by building in memory a predictive model of the world. The paper complements a 1988 report on the Organization of an Autonomous Learning System, henceforth referred to as the “Target Article” (Kanerva 1988a).
Neural-net associative memories for high-dimensional (aka “wide”) vectors were studied already in the 1970s, but operations of the right kind for computing with vectors were found only in the 1990s, starting with Plate’s Holographic Reduced Representation (Plate 1991). Computing with vectors is subtle in that it takes place in superposition.
We compare computing with vectors to conventional computing with numbers and argue for modeling and simulating human and animal cognition with a von Neumann-like architecture for computing with vectors. We begin with an overview of conventional computing and identify corresponding circuits for computing with vectors. We then discuss reasons to distinguish between working memory and long-term memory. The vector math is carried out in the working memory and the results are stored in the long-term memory and ultimately drive action. We make no attempt to map the working memory to neural circuits in the brain, but observe that the circuits of the cerebellum are laid out beautifully in 3D for a massive long-term memory for wide vectors (see e.g. Eccles et al. (1967) and figures 7.2 and 7.3 of Llinás et al. (2004)): the 3D layout resembles the wiring of magnetic-core memories of the 1960s, with their bit planes corresponding to the flat dendritic trees of Purkinje cells. Kanerva (1993) illustrates step by step the cerebellum’s likeness to computer memory and to ANNs. The memory circuit consists of mossy fibers, granule cells, Purkinje cells, and climbing fibers. The axons of granule cells bifurcate into parallel fibers that run perpendicular to Purkinje cells’ flat dendritic trees.
The rest of the paper and the target article are about applying the architecture to autonomous learning by robots, modeled after learning by humans and animals with sensorimotor systems controlled by advanced brains. The paper represents a systems approach that is aligned with James Albus’s ideas about robots with human-like faculties. His book on Brains, Behavior, and Robotics (Albus 1981) came out a decade before ideas about computing with vectors in superposition began to take hold, but is still relevant and worth reading.
3 Traditional Computing with Numbers
This section serves readers who know computers from having programmed them in high-level languages but know less about the hardware within. The lesson is that the computer in itself is exceedingly simple and that its success comes from programming.
The traditional computing architecture, also known as the von Neumann architecture, consists of a memory for storing programs and data (Random-Access Memory, RAM), and a processor that steps through the program and does the math (Central Processing Unit, CPU). The memory is made of sequentially numbered storage locations, and a location’s number serves as its address. Even small computers have memories with millions of storage locations.
A storage location is a piece of hardware that can hold a small amount of data, for example a number, a letter, or a simple command for the processor to carry out, such as adding two numbers. A storage location can be visualized as a fixed number of bits, 32 for example, and the programmer decides what the bits in any particular location—at any particular address—mean.
The CPU has two functions, step-by-step running of the program and the math that is built into the computer’s circuits. The math usually includes addition, subtraction, multiplication, and division of numbers, and logic operations on strings of bits (AND, OR, NOT, shift and rotate). More complicated operations are made with programs that combine these basic operations. For example, a program to calculate the mean of a set of numbers copies the numbers, one at a time, from the memory into the processor, adds it to the sum, divides the final sum by the count, and stores the result in memory for later use—for comparing the mean heights of two groups of people, for example.
3.1 Random-Access Memory—RAM
Random Access is an all-important property of the memory. It means that the contents of any storage location are immediately available to the processor. The processor merely needs to “know” the location’s address. Random access is in contrast to storing data on tape or disk that make the processor wait for the addressed memory location to be under the read/write head. The “miracle” of random access is achieved with an address-decoder circuit.
3.2 Linked Lists
Random access allows lists of numbers to be stored so that any number in the list can start the retrieval of the rest of the list. The numbers in the list are used as addresses of storage locations and as data to store, and the list is stored by stepping through it one number at a time, and using it as the address in which to store the next number. Linked lists (also called pointer chains) are standard fare in list-processing languages such as Lisp, and provide a way to retrieve a list starting with any number in it. When vectors are linked in a list like that, they can become a predictive model of the world, as discussed under “predicting” in subsection 6.3.
Finally, computers get their inputs from the world and deliver their outputs through transducers that convert between physical signals and their representation in bits.
4 Computing with Vectors
By a $D$-dimensional vector we mean a series of $D$ numbers treated as a coherent unit. The traditional (von Neumann) architecture for computing with numbers is readily extended to computing with wide vectors (e.g., $D$ = 10,000). The “new” architecture is essentially the old: central processing unit (CPU) to carry out the vector math, Random-Access Memory (RAM) for wide vectors, and circuits for input and output. A traditional computer would control the running of the program, with a vector-arithmetic unit running as a coprocessor, for the time being at least.
4.1 Experimenting with Vectors
We wrote a program for identifying languages using four operations on vectors: addition, multiplication, permutation, and dot product (Joshi et al. 2017). The experiment was motivated by people’s ability to identify languages by how they sound, without knowing any of the words. Rather than speech, we used about a million bytes of text for each of 21 European Union languages, transcribed in Latin alphabet of 26 letters and a symbol for the space between words. The challenge was to identify languages by how they look in print, without knowing any of the words. We made a 10,000-dimensional (10K) profile vector for each language and for each test sentence, and identified a test sentence with the language that had the most similar profile.
An analysis of the language profiles clustered them in four families: Slavic, Baltic, Romance and Germanic, with English between Romance and Germanic. Of 21,000 test sentences, 97.3% were identified correctly, and mistaken ones fell mostly in the same language family. The experiment was ran on a laptop computer in a single pass over the data and the test sentences, and took less than 8 minutes.
Profiles for languages and test sentences were computed with the same algorithm, and were compared to each other with the cosine. We started with 10K vectors for letters, computed 10K vectors for trigrams, and added them into 10K vectors for profiles. Each of the 27 letters was assigned a 10K independent random seed vector of 1s and 1-1s. The trigrams of the data are all its three-letter sequences and they were encoded into 10K trigram vectors with permutation and multiplication: permute the first letter (its 10K code) twice, permute the second letter once, take the third as is, and multiply the three. Permutation was done by rotating vector coordinates and multiplication was done coordinate by coordinate. Finally, a 10K profile vector was computed by summing over all the trigram vectors of the text: the profile and test vectors were sums of products of permutations of seed vectors for letters. The meanings of “add,” “multiply,” “permute,” and “similar” are explained next.
4.2 Instruction Set for Vectors
What operations would make for useful computing with vectors? Linear algebra perhaps, a well-established branch of mathematics that has a major role in artificial neural nets. However, we want operations that do not change dimensionality, so that an output of an operation can serve as an input to further operations, and also as an address to memory. However, the product of two vectors in linear algebra is either a matrix or a number, and so we need a different way to multiply vectors. We do it coordinatewise, known as the Hadamard product.
Three simple operations make a surprisingly powerful system to compute with: vector addition, coordinatewise multiplication, and the shuffling of vector coordinates (permutation). The vectors can be binary or integer or real or complex—the computing power comes more from high dimensionality (e.g., $D$ = 10,000) and randomness than from the nature of vector components.
The operations have been discussed at length in the literature (Plate 2003; Kanerva 2009, 2023) and can be illustrated with binary vectors (equivalent to vectors of 1s and 1-1s). Addition is by coordinatewise majority rule with ties broken at random (equivalently, ordinary vector addition followed by applying a threshold to each coordinate); multiplication is coordinatewise Exclusive-Or (XOR); and permutations reorder vector coordinates. The algebra of addition and multiplication resembles ordinary arithmetic in regard to associativity, commutativity, invertibility, and distributivity. Permutations have no counterpart in ordinary arithmetic. They are invertible, most do not commute, and they distribute over both vector addition and multiplication. Permutations of coordinates add all finite groups up to size $D$ into the vector math.
The operations allow traditional data structures (sets, sequences, lists, trees, stacks, etc.) to be represented and operated on in superposed vectors (Kleyko et al. 2022). For example, sets can be encoded with addition, the binding of variables to values with multiplication, sequences with permutations, and data structures at large with combinations of the three operations. For a simple example, a vector that means $x=a$ can be encoded by the product P=XAP=X*A where $X$ and $A$ are $D$-dimensional vectors representing the variable $x$ and the value $a$, and * is coordinatewise multiplication (XOR for binary vectors). The three operations are also used to decode a composed vector into its constituent vectors: the value of $x$ in $P$ can be recovered by multiplying $P$ with (the inverse of) $X$, namely XP=X(XA)=($X$)A=AX*P=X*(X*A)=(X*X)*A=A, and by noting that XOR is its own inverse so that XXX*X evaluates to a vector of 0s. In ordinary computing, the variable $x$ is represented by an address of a memory location and is buried in the program code, and the value $a$ is represented by the contents of the location. The three operations also allow regularities in data to be captured in vectors, opening a path to statistical learning.
A fourth operation, based on the dot product, measures the similarity between vectors (Hamming similarity, cosine, or Pearson correlation for example). The idea is that when vectors are similar their meanings are similar. Nearly all pairs of vectors in a high-dimensional space are dissimilar (effectively orthogonal)—randomly drawn pairs of vectors are approximately orthogonal. An endless supply of approximately orthogonal vectors is a major reason for insisting on high dimensionality.
4.3 Computing in Superposition
Computing in superposition goes by many names, depending mainly on the intended audience—whether mathematicians, physicist, cognitive scientists, linguists, engineers, or computer scientists: Holographic Reduced Representation (HRR), Vector Symbolic Architecture (VSA), Context-Dependent Thinning, Hyperdimensional Computing, and Semantic Pointer Architecture (SPA). Each item or element or feature or thing or piece of data—each vector—that is included in a superposed vector is distributed evenly over all vector components—any subset of vector components represents the same thing as the entire vector, only less precisely. Representation of this kind is called “holographic” or “holistic,” as opposed to the conventional localist representation where a variable is represented by an address of a memory location and its value by the contents of the location. The encoding of $x=a$ with the vector $P$ in subsection 4.2 is an example of holistic representation: P=XAP=X*A.
Computing in superposition begins with $D$-dimensional seed vectors chosen at random (with i.i.d. components), to represent basic entities such as variables and names. In essence, they are symbols or tokens to compute with. They can be used as addresses to memory and as data stored in memory, and as inputs to the math operations that combine them into new vectors, which in turn can address the memory, be stored as data, and serve as inputs to the vector operations. New representations are made from existing ones with explicit calculation, which is fundamentally different from the generation of representations in an autoencoder or in the layers of a deep neural net as it is trained. Distributivity and invertibility of operations also mean that codes for composite objects can be decoded into their constituents, in some cases at least.
The operations allow a high degree of parallelism. Because addition and multiplication happen coordinatewise, independently of other coordinates, all $D$ (= 10,000) can be done at the same time. Furthermore, because the operations are simple, it is feasible for each dimension to have its own processor circuit. Computing with vectors in superposition can then be seen as each coordinate having its own simple computer, all computing the same thing but at exceedingly low precision—down to 1-bit. On the aggregate, however, they compute with high precision if the dimensionality is high enough—and even if some of the simple circuits malfunction! In contrast, traditional circuits for computing with numbers are complicated and are expected to work flawlessly. The simplicity of the coordinatewise operations makes it possible to distribute them throughout the system, reducing the need to move data before it can be computed on. This is called “in-memory computing.”
5 Ideas for Architecture from Biology and Psychology
How does computing with vectors fare in the light of biology and behavioral sciences? Psychological research in the 1960s and ’70s identified two levels of organization of human memory, short-term working memory and long-term data store (Atkinson & Shiffrin 1968; Anderson & Bower 1973); see also Wikipedia article on “Atkinson–Shiffrin memory model”). Representations are assembled and analyzed in the working memory and stored in the long-term memory (Tulving & Donaldson 1972; Tulving & Thomson 1973). These early models summarize behavior without accounting for underlying mechanisms and their math.
The psychologists’ models are similar to traditional computing that assembles and interprets data structures in the central processing unit (CPU) and stores them in the random-access memory (RAM). Computing with vectors follows the same organization: vectors are encoded and decoded in the working memory (counterpart of CPU) and stored indefinitely in the long-term memory (counterpart of RAM). We will discuss the long-term memory first.
5.1 Long-Term Memory for Vectors
Memory allows past experience to bear on the present situation. Associative memory, also called content-addressable (these terms lack a precise definition), has been an object of psychological inquiry for over a century (James 1890; Raaijmakers & Shiffrin 1981), and mathematical models have been proposed since the early days of artificial neural nets (see Hinton & Anderson 1981, for a review). Sparse Distributed Memory (Kanerva 1988b) takes the idea a step further by assuming that the memory could have billions of locations and that its function is long-term data storage. In the rest of this paper, “long-term memory,” “associative memory,” “high-dimensional RAM,” “Sparse Distributed Memory,” and “SDM” mean the same thing and are often referred to simply as the memory. Its similarity to computer memory is significant: the computer RAM stores numbers and is addressed by numbers, an associative memory such as the SDM stores vectors and is addressed by (cued with) vectors.
Psychologically realistic models of memory for vectors, such as the Sparse Distributed Memory (Kanerva 1988b, 1993), consists of locations for storing $D$-dimensional vectors. Each location is associated with a $D$-dimensional address. A location is activated when the memory is probed/cued/addressed with a vector that is similar to the location’s address. The similarity criterion is chosen so that multiple locations are activated, but only a tiny fraction of all possible locations (Kanerva 1993). An activated location either accepts a $D$-dimensional vector (write operation) or outputs a $D$-dimensional vector (read operation). The final output of a memory read is computed from the outputs of all activated locations. Although the number of possible locations borders the incalculable, it is possible to simulate random access by distributing data over multiple locations and reconstructing it upon retrieval.
There is a difference of lesser importance to note: storing a number in a computer memory replaces the old contents of the activated/addressed location; storing a vector in an associative memory adds it to the contents of all activated locations—the newly stored data are superposed.
An associative memory’s likeness to the cortex of the cerebellum is remarkable (Marr 1969; Albus 1971; Kanerva 1993), and the numbers agree with the need for a high-capacity memory. With 50 billion granule cells (Llinás et al. 2004) representing locations for storing of vectors, the human brain could store one vector every minute (more than 10 million bits a day) over a life span of a 100 years.
Catastrophic forgetting is a known phenomenon of artificial neural nets. It is in part due to the limit on the amount of information that any given “hardware” can store, and in part to training a monolithic network with error back-propagation (French 1999). Catastrophic means that when the capacity limit is reached, the net begins to produce outputs that are unrelated to the vectors it was trained on—it’s like a phase transition in physics where the properties of a substance change suddenly, as when super-cooled water turns into a block of ice. The alternative is graceful degradation, which can be achieved with a memory that is organized in storage locations whose contents can change but addresses are fixed, as they are in ordinary computer memory, in the Sparse Distributed Memory, and apparently also in the cortex of the cerebellum. Parallel-fiber synapses with Purkinje-cell dendrites are modifiable and make up over half the brain’s synapses; (Llinás et al. 2004). However, a small fraction of synapses (1%) should be fixed according to the theory, namely, synapses between mossy fibers and granule cells that determine which locations/granule cells to activate. They act as an address decoder and provide a framework for recording a lifetime of experience.
5.2 Short-Term Working Memory
There is more to consider when computing with vectors in superposition: just how many vectors to encode into a single vector? The number is limited by vector dimension and determines the amount of structure that a single vector can hold (Gallant & Okaywe 2013; Frady et al. 2018), beyond which it is necessary to store vectors in the long-term memory. Psychologists call it chunking.
The human mind can recall large amounts of data that has been organized in chunks (and in chunks of chunks, and so on) and stored in memory. George Miller’s “Magical Number Seven, Plus or Minus Two” is the classic on the topic (Miller 1956). The working memory combines pieces of data into chunks and stores them in the associative memory, and recalls chunks into the working memory, to compose them into further chunks or to break them down into their constituent parts.
Miller’s magical number suggests that a chunk should combine fewer than ten items. That is also about the number of vectors that can be encoded into a single vector and readily decoded with the vector math. The number grows nearly linearly with vector dimensionality $D$ (Frady et al. 2018). With psychology and the vector math as clues, we assume that a 10,000-dimensional working memory can hold and operate on up to ten vectors at a time.
6 Modeling the World with Vectors
To help us understand intelligence of animals (and of ourselves), we liken them (and us) to robots whose behavior is controlled by “brains” that compute with wide vectors. Albus (1981) makes a case for humanlike robots, and Kanerva (1988b), Plate (1991) and Piantadosi et al. (2024) argue for representing concepts with high-dimensional vectors. The robot has sensors and actuators for interacting with the world, an associative memory for storing a model of the world, and a working memory where the world model is constructed and interpreted with the vector math (Teeters et al. 2023).
6.1 The Focus
The contents of the working memory are summarized in a single $D$-dimensional vector called the focus. The idea of a focus at the core of a cognitive computing architecture goes back to 1988 (Kanerva 1988a, b) but has not been adopted widely. However, it deserves a new look now that von Neumann-like computing with vectors in superposition has become understood Plate (1991). The sensors feed into the focus, the activators are driven from the focus, the memory is addressed by the focus, and vectors in and out of the memory pass through the focus. The vector in the focus is the system’s (robot’s) state and it represents the robot’s subjective experience at the moment—the focus is the robot’s subjective self (an objective state would have to include the contents of memory as well).
The focus is a theoretical construct, meant to capture the idea of an integrated self. For example, any of a number of things can bring to mind a specific experience from the past; the trigger can be a specific sight or sound or smell or ache or pain, or some combination of them. Thus any one of them can serve as an address to memory, to evoke the past experience. Similarly, we recognize a person by how they look or sound or move or act: all those different modalities refer to—and retrieve from memory—the same person.
It seems befitting that the focus is limited to a single vector because our minds can attend to only a few things at a time (cf. Miller’s magical number 7). Limited capacity is behaviorally realistic. Furthermore, because the focus is at the crossroads of sensors, actuators, and memory, an experience created by the world through the senses can also be created from the memory. In computing terms, they reside in a common $D$-dimensional space, allowing the history to be stored as a linked list; see subsection 3.2.
As an added benefit, representing sensory modalities and action in common mathematical space allows generic algorithms to be used for learning. It also means that we need to turn physical (and chemical) signals with very different properties (sight, sound, taste, smell, temperature, pressure, etc.) into $D$-dimensional vectors for the working memory. The best ways to do it will be worked out over time. Meanwhile, standard signal-processing techniques together with random projections have been used successfully. For example, letters of the alphabet have been represented by random vectors for written-language identification (Joshi et al. 2017), words of a language have been represented by random vectors for comparing word meaning (Kanerva et al. 2000), and Fourier coefficients for speech have been randomly projected to a $D$-dimensional space for speaker identification (Huang et al. 2022).
Looking ahead, we need to learn how nervous systems and brains prepare sensory input for further processing, and build that into artificial systems. For example, the cochlea of the inner ear analyses sound into frequencies—it Fourier-transforms the sound before passing it on to the rest of the brain. Regarding vision, the optic nerve brings in information along about 1.2 million axons per hemisphere (Sterling & Demb 2004) and the primary visual cortex distributes it among some 140 million neurons (Leuba & Kraftsik 1994)—a 100-fold increase.
6.2 The Passage of Time
To act intelligently in the world, the actor—an individual or a robot—must deal with the passage of time. We are concerned here with subjective time and represent it with a sequence of moments, and a moment with the focus vector. A system’s history is then a life-long sequence of vectors for moments of time. The history can be divided into episodes defined by shared context. The challenge is to store the history in a way that allows a system to learn from experience.
Short sequences, up to ten or so, can be encoded into a single vector in the working memory and treated as a chunk. Joshi et al. (2017) encoded short sequences of letters ($N$-grams) as chunks and used them to identify languages. The long-term associative memory makes it possible to deal with longer sequences by storing the history in the long-term memory as a life-long linked list; see subsection 3.2. The vector for the present moment—the present state, the focus—can then be used as an address to trigger the recall of similar past episodes.
6.3 Sensing and Acting
Sensing. Animals have elaborate sensory systems that receive data about the world: the brain “swims” in a torrent of data and needs to find regularities and patterns in the data stream. A wealth of sensory input comes from the body itself. In fact, the body can be regarded as a part of the world from which we receive input and on which we act.
Detecting. Recognizing previously encountered states makes it possible to detect irregularities and anomalies that can serve as an alarm, for example. It does not take much to see that “wrord” is anomalous, for example.
Detection can be accomplished with very basic learning: store the vector for each moment in the memory using the vector itself as the address—this is called “autoassociation.” Frequently occurring similar vectors are reinforced and infrequent vectors become background noise. When the sensory data for a moment arrives, it can be used to address the memory to determine whether it represents something new or something “seen” before. If it recalls a dissimilar vector the new vector is stored, if the two are essentially the same the new vector can be ignored, but if they are similar, the difference between the two can be analyzed with the vector math to choose an action to take.
Predicting. Intelligent behavior requires more than the ability to detect whether I have seen this before or been here before. We need to look ahead: what to expect next and how to act accordingly. We need to exploit regularities in data that unfold over time.
Prediction can be accomplished by storing the system’s history in the memory as a linked list, as described in subsection 3.2. When a vector for a moment recurs and is used to address the memory, it recalls—and predicts—the next moment. A moment later the prediction can be compared to what actually happens—this is essentially anomaly detection that can initiate further action.
Acting. To talk of acting implies the existence of an actor. We liken it to a robot or an animal or a human being coping with a seemingly boundless and indifferent world. To exist in the world, the actor needs to acquire resources from the world (material and energy) and to steer out of harm’s way—to eat and not to be eaten. The idea of an actor implies the notion of “self,” an individual that is separate from the rest of the world—we use the word “individual” for actors of different kind but not for the world. Individuals/actors have some degree of autonomy. The individual’s actions become a part of the world model.
Reacting. Reaction is a response to a stimulus. Some animal behaviors are passed on genetically and other are learned during life. The model deals with both—see discussion of the preference function in the target article. Learning in real time takes advantage of the memory’s ability to predict. The actions and reactions of today’s robots are mostly programmed, and very little learning takes place in real time out in the field.
Interacting implies two or more parties reacting to each other’s actions. If one of the parties is the indifferent world, the individual can act on it, to discover how the world works. When the interaction is between two individuals, each one can learn about the other by probing, but there is more: rather than being indifferent, the two can cooperate or compete, in which case my model of the world needs to include a model of you, which needs to include a model of me, and so on. To avoid infinite regress, we project ourselves to the other individual—act as if the other were like ourselves—and regress only a step or two beyond that.
What is the individual’s relation to its model of the world: how is the model used? We can think of the model as a database that the individual consults, to evaluate alternative actions and to choose one that is appropriate for the present situation. Most of today’s robots work like that, but it is not what actually happens in brains because it assumes a high-level executive (a homunculus) that runs the show. The use of the model by the individual needs to be more organic. Looked at in that way, rather than an individual having a model of the world, the individual is the model (and the model is the individual).
Interaction with the world and its individuals are the data for the world model. An individual’s interaction is experienced as a sequence of states of the focus that include both sensory and motor components. When stored in memory as a linked list, it allows the individual to predict and to act according to the prediction. The details are worked out in the target article (Kanerva 1988a) and will not be repeated here. We merely comment on it.
7 Autonomous Learning (Kanerva 1988a)
in Retrospect
Computing with wide vectors tries to capture the organization of animals that learn and adapt. It is based on the idea that an associative memory can learn predictive modeling of the world with a circuit that is strikingly similar to the cerebellum’s. In psychologists’ terms, the associative memory is a long-term data store.
The architecture includes a working memory. Its contents are summarized in a single vector called the focus. The state of the focus represents the system’s/robot’s/individual’s subjective experience from moment to moment.
The 1988 paper has nothing concrete to suggest about how the working memory actually works. The paper merely refers to an “Encoding Problem” as yet to be solved. The gist of the encoding problem was solved by Plate by computing in superposition with Holographic Reduced Representation Plate (1991). Other than that, the architecture is conventional (von Neumann) and suitable for both symbolic (Gayler 2003) and statistical computing (Joshi et al. 2017).
The 1988 paper refers to varying the amount of information that a given data source or destination contributes to the focus vector, but does not tell us how actually to do it. Superposition suggests an answer. When each data source and destination is encoded into a $D$-dimensional vector, the vectors can be weighted by importance before adding them to the focus. For example, we can prioritize sight over sound, or vice versa, depending on the context. We would no longer refer to feature vectors, as that suggests vectors being partitioned into fields for features. Every “feature” that is encoded into a vector is present in every coordinate. In fact, what we traditionally think of as features are more like afterthoughts, to help us compartmentalize the world. If everything in the world were green, color as a feature would not exist.
8 What Next?
Computing with wide vectors seems particularly suited for simulating—and possibly understanding—intelligence of animals with advanced sensorimotor systems and brains (Albus 1981; Kanerva 1988b; Piantadosi et al. 2024). Much needs to be worked out, however, to develop the idea into a viable technology. Fortunately, developing and testing does not require excessive hardware or computing. Ample opportunity to experiment, innovate and invent is within reach (see Yang 2023, for an example). Here are some of the challenges we face.
Programming and Control. Whether and how to do the actual programming with superposed vectors is unclear. For now, conventional programs running in conventional computers control the operation of the hyperdimensional computer in simulation or prototype. However, people are beginning to think also about vector-based control structures (Tomkins-Flanagan & Kelly 2024, e.g.).
Sparse Vectors. Matching sets of operations for dense vectors have been known from the start (1990s). They are easy to compute with but inefficient. We need good sets of operations for sparse vectors without compromising computing power. The algebra of the operations must allow multiple vectors to be encoded into a single vector, operated on, and the result decoded into known vectors.
Associative Memory. Computing with vectors depends crucially on a high-capacity associative memory for wide vectors (a high-dimensional RAM), that does a fast near-neighbor search in a huge and sparsely populated address space. When vectors are stored in an associative memory with themselves a addresses (called autoassociative) they can become fixed points with basins of attraction of a dynamical system, allowing rapid storing and retrieving of vectors. The Sparse Distributed Memory (Kanerva 1988b) demonstrates the concept but is inefficient. The circuits of the cerebellum can suggest more efficient designs, for example by judicious use of randomness and sparseness (Jaeckel 1989).
Noise. One of nature’s marvels is how we recognize a person or an object or a place or a situation—anything—as being the same now as at some earlier time, and distinct from other things, in spite of the data arriving at the senses never repeating exactly. The tendency to categorize things based on what we expect of them, seems to be built in. But even “built in” needs explaining: a physical process that sustains the behavior. Without going into it here, we merely presume that high dimensionality plays an important role, and instead comment on how variability in input is handled in cognitive modeling today.
Objects, etc., are treated as points in a feature space that has a similarity measure. Prototypical objects occupy regions of the space, and nearby points are considered to be noisy versions of them. This works when most points are far away, as they are in a high-dimensional space. If the prototypes are stored in memory as fixed point, a noisy version that falls in a basin of attraction is recognized with a slight delay typical of attractor networks.
Another source of “noise” cannot be ignored. It has to do with encoding and decoding of superposed vectors. For binary vectors, multiplication and permutation are invertible and preserve information, but addition is only approximately invertible, and so the decoding of a vector sum puts out noisy vectors. That limits the number of vectors included in a sum, and calls for cleaning up of vectors decoded from a sum before further operating on them.
Neuroscience. We have a reasonable idea of how a long-term memory in the brain is organized (Albus 1971; Kanerva 1993; Llinás et al. 2004; Jaeckel 1989, cf. the cerebellum:) but none for the working memory. If computing with vectors produces human-like or animal-like behavior with energy comparable to brain’s, we can hope to understand and identify brain’s circuits for working memory.
Psychology and Psychophysics. Psychologists have probed human (and animal) behavior in countless experiments, and conjectured about processes that underlie memory recall, response time, attention span, interference, ability to discriminate, and so on. Relating the observations to computing with wide vectors can deepen our understanding of the underlying processes and lead to the engineering of more lifelike robots, for example.
Modeling of brains with wide vectors also opens new vistas in experimental psychology. As the theory and models improve, they will suggest experiments to confirm or counter hypothesized underlying mechanisms. This can benefit both psychology and engineering.
Language and logic are the apex of symbolic behavior and have even been considered as defining intelligence. Symbolic processing also formed the early core of AI—until we tried to build systems that learn by interacting with the world. Where logic and language failed, artificial neural nets were thought to succeed, leading to today’s deep neural nets and large language models. Impressive as they are, their energy requirements so exceed the efficiency of brains, as to rule them out as models of how brains work and make us intelligent.
Human and animal intelligence is based on both learning from sensor data—exploiting regularities in the statistics—and reasoning, even in the absence of fully developed language. Artificially intelligent systems need to be capable of both, which is what computing with vectors aims at. The algebra of the operations makes symbolic processing possible (e.g., the explicit encoding, storage and decoding of structured data), and high-dimensional representation can include probabilities without explicit tallying and bookkeeping.
9 Summary
The paper outlines a traditional (von Neumann-like) architecture for computing with wide vectors and argues for its use for autonomous learning in real time by systems such as robots. It is also meant to help us understand how humans and animals learn. It combines ideas from psychology (short-term working memory and long-term data store, chunking), neuroscience (cerebellum-like circuit for an associative memory), computer science and engineering (central processing unit, random-access memory, sequence as a linked list), and mathematics (algebra of vector operations, effective orthogonality of random vectors). Peculiar about computing with vectors is that it takes place in superposition. Variables and values are vectors of a common mathematical space, and so is a variable having a certain value. The memory for storing the vectors is also addressed by the vectors—it’s a “high-dimensional RAM.”
The subjective state of the system is defined by a single vector called the focus. It summarizes the contents of the working memory. The sensors feed into the focus, the actuators are driven from the focus, and the memory is accessed through the focus. The passage of time is represented by a sequence of vectors in the focus. When stored in the memory as a linked list, it becomes a predictive model of the world.
Autonomous learning requires that some sensory states start out as inherently desirable (satisfy a need) and some as inherently undesirable (painful). A model of the world that includes the system’s own actions can then be used to learn sequences of actions that lead to desirable states and avoid harmful ones.
The architecture of the long-term memory (Sparse Distributed Memory) is shallow. It has one fixed layer for activating memory locations and one variable layer for storing and recalling data in the activated locations. The fixed layer corresponds to the address-decoder circuit of a computer memory (RAM), and the variable layer corresponds to memory stacks that store the data. This architecture is suggested by the wiring of the cortex of the cerebellum, which contains over half the brain’s neurons and synapses. The variable layer can be trained in real time by the perceptron learning rule, and learning can happen in only a few rehearsals.
The vector math is carried out in the working memory (counterpart of the Central Processing Unit, CPU) with an instruction set consisting of coordinatewise addition and multiplication, and the permutation of coordinates. Similarity of vectors is based on the dot product. The operations can combine symbolic processing and statistical learning.
Interaction with the world happens through sensors and actuators. Input is made into wide vectors for the working memory/focus with sensor-specific pre-processing, and the focus is decoded for commands to actuators with motor-specific post-processing. Pre- and post-processing can be engineered, inherited or learned early in the system’s existence in multiple passes with backpropagation, and then fixed.
Computing with vectors addresses many of the same issues as deep neural nets but with an architecture more like that of a traditional computer, or of animals with advanced brains. It should do well at simulating human and animal cognition, but that is for future research to determine.
Supplemental Material
RIACS Report 88.14 on the Organization of an Autonomous Learning System (Kanerva 1988a), referred to here as the Target Article.
Acknowledgments
This work was supported in part by CoCoSys, one of the seven centers sponsored by the Semiconductor Research Corporation (SRC) and DARPA under the Joint University Microelectronics Program 2.0 (JUMP 2.0); and in part by the U.S. National Science Foundation (NSF) under Cooperative Agreement No. 2433429, “NSF Al Research Institute on Interaction for Al Assistants (ARIA).”
Sincere thanks to professor Peter Denning for discussions that lead to writing of the paper, and to an anonymous reviewer for comments that eventually lead to its publication. Special thanks to professor Bruno Olshausen and his Redwood Center at UC Berkeley, for creating an environment favorable for scientific research and discovery.
Publisher’s Note. All claims expressed in this article are solely those of the author and do not necessarily represent those of their affiliated organizations, or those of the publisher, the editors and the reviewers. Any product that may be evaluated in this article, or claim that may be made by its manufacturer, is not guaranteed or endorsed by the publisher.